OpenStudy (anonymous):

What is 14 10/52 x 18 2/46?

OpenStudy (anonymous):

\[14*\frac{10}{52}*18\frac{2}{46}\] \[\frac{140}{52}*\frac{36}{46}\] \[\frac{140*36}{52*46}\] \[\frac{5040}{2392}\] 2,10702341